Apply your knowledge gaining from this course in describing how you convert a function into a relation and a relation into a function?

1
Expert's answer
2021-05-19T18:11:27-0400

If we have a function "f: X\to Y", then define a relation "R\subset X\times Y" in the following way: "R=\{(x,y)\ |\ y=f(x) \}\subset X\times Y."


On the other hand, for a functional relation "R\subset X\times Y" let us define a function "f: X\to Y"

in the following way: "y=f(x)" if and only if "(x,y)\in R."